solve for b a= 1/2bc c is squared
i tried 2a/c squared

Answer by sdmmadam@yahoo.com(530)

a= 1/(2bc^2)----(1)
Multiplying by (2bc^2)
aX(2bc^2)= 1
2abc^2 = 1
Dividing by 2ac^2
(2abc^2)/(2ac^2) =1/(2ac^2)
b = 1/(2ac^2)
Answer: b = 1/(2ac^2)
Verification: Putting b = 1/(2ac^2)in (1)
RHS = 1/[2X[1/(2ac^2)] Xc^2]
=1/[2c^2/2ac^2]
=1/[(1/a)] (cancelling 2c^2)
= a = LHS
Note: Instead of going through the steps methodically
you may multiply (1) by b and divide by a to get immediately b = 1/(2ac^2)
which is the easiest way of handling the problem